OpenStudy (anonymous):

Suppose a particular bank gives 3% interest on its savings accounts, compounded monthly. If one opens a savings account and deposits $ 300 each month for a total of 11 years, how much money will be in the account after the last deposit is made?

OpenStudy (anonymous):

\[-\frac{D \left(1-(1+R)^N\right)}{R}\]

OpenStudy (anonymous):

D=300 R= .3/(100*12) n=11*12

OpenStudy (anonymous):

I plugged in those numbers into the equation and got 0 which is incorrect